### v2

# SASSL: Enhancing Self-Supervised Learning via Neural Style Transfer

Abstract
Self-supervised learning relies heavily on data augmentation to extract meaningful representations from unlabeled images. While existing state-of-the-art augmentation pipelines incorporate a wide range of primitive transformations, these often disregard natural image structure. Thus, augmented samples can exhibit degraded semantic information and low stylistic diversity, affecting downstream performance of self-supervised representations. To overcome this, we propose SASSL: Style Augmentations for Self Supervised Learning, a novel augmentation technique based on Neural Style Transfer. The method decouples semantic and stylistic attributes in images and applies transformations exclusively to the style while preserving content, generating diverse augmented samples that better retain their semantic properties. Experimental results show our technique achieves a top-1 classification performance improvement of more than $2\%$ on ImageNet compared to the well-established MoCo v2. We also measure transfer learning performance across five diverse datasets, observing significant improvements of up to $3.75\%$. Our experiments indicate that decoupling style from content information and transferring style across datasets to diversify augmentations can significantly improve downstream performance of self-supervised representations.
1 Introduction
Data labelling is a challenging and expensive process, which often serves as a barrier to build machine learning models to solve real-world problems. Self-supervised learning (SSL) is an emerging machine learning paradigm that helps to alleviate the challenges of data labelling, by using large corpora of unlabeled data to pre-train models to learn robust and general representations. These representations can be efficiently transferred to downstream tasks, resulting in performant models which can be constructed without access to large pools of labeled data. SSL methods have shown promising results in recent years, matching and in some cases exceeding the performance of bespoke supervised models with small amounts of labelled data.
Given the lack of labels, SSL relies on pretext tasks, i.e., predefined tasks where pseudo-labels can be generated. Some examples include contrastive learning (Chen et al., 2020a; He et al., 2020), clustering (Caron et al., 2021; 2020; Assran et al., 2022), and generative modeling (He et al., 2022; Devlin et al., 2018). Many of these pretext tasks involve training the model to distinguish between different views of the same input and inputs corresponding to different samples. For these tasks, the way input data is augmented is crucial for the network to learn useful invariances and extract robust representations (Chen et al., 2020a). While state-of-the-art augmentations incorporate a wide range of primitive color, spectral and spatial transformations, they often disregard the natural structure of an image. As a result, SSL pre-training methods may generate augmented samples with degraded semantic information, and may be less able to capture diverse visual attributes.
To overcome this limitation, we propose Style Augmentations for Self Supervised Learning (SASSL), a novel SSL data augmentation technique based on Neural Style Transfer to obtain stronger augmented samples via semantically-aware preprocessing. In contrast to handcrafted augmentation techniques operating on specific formats (e.g. pixel or spectral domain), SASSL disentangles an image into perceptual (style) and semantic (content) representations that are learned from data. By applying transformations exclusively to the style of an image while preserving its content, we can generate augmented samples with diverse styles that retain the original semantic properties.
Our contributions:
  • We propose SASSL, a novel data augmentation technique based on Neural Style Transfer that naturally preserves semantic information while varying stylistic information (Section 4).
  • We empirically demonstrate improved downstream task performance by incorporating our method into MoCo v2, without additional hyperparameter tuning. We report more than $2\%$ top-1 accuracy improvement on the challenging ImageNet dataset (Section 5.1).
  • We show our augmentation method learns more robust representations by measuring their transfer learning capabilities on five diverse datasets. SASSL improves linear probing performance by up to $3.75\%$ and fine-tuning by up to $1\%$ on out-of-distribution tasks (Section 5.2).
  • We observe that balancing stylization using representation blending and image interpolation performs best, and adding external style datasets for transfer can improve performance (Section 5.3).
(a) Proposed data augmentation pipeline.
(b) Style transfer preprocessing.
Figure 1: Towards diverse SSL data augmentation via Neural Style Transfer. We propose SASSL, a novel augmentation method based on Style Transfer that generates semantic-aware transformations of pre-training images by operating exclusively on their style. Given a pre-training image and an external style representation, SASSL combines the content attributes of the image with the external style, producing an augmented view that better preserves semantic attributes. By controlling the stylization strength via feature and pixel interpolations, we empirically show SASSL promotes robust representations by incorporating Style Transfer to default SSL augmentation pipelines.
2 Related Work
2.1 Data Augmentation in SSL
Typical data augmentation methods in SSL applied to vision tasks correspond to image cropping and resizing, flipping, rotation, color augmentation, noise addition, and solarization. Examples of methods using these are MoCo (He et al., 2020), SimCLR (Chen et al., 2020a), BYOL (Grill et al., 2020), and SimSiam (Chen & He, 2021), among others. Other research has explored how to select augmentation hyperparameters to learn more robust and general features (Wagner et al., 2022; Reed et al., 2021; Tian et al., 2020). In contrast, our method proposes a new preprocessing technique that can be incorporated in any existing augmentation pipeline, obtaining improved performance without additional hyperparameter tuning.
Novel augmentation techniques have been proposed for both supervised and self-supervised learning. For example, Han et al. (2022) propose a novel method to apply multiple data augmentation techniques in a local fashion by splitting the image into patches. This is aligned with other local augmentation techniques such as random erasing (Zhong et al., 2020), CutMix (Yun et al., 2019), Mixup (Zhang et al., 2017), among others. While these methods focus on masking or severely distorting image patches, disregarding their semantic information, our approach is a global transformation that decouples an image into its appearance and semantic attributes. This allows applying transformations exclusively to its appearance without affecting the image semantics.
2.2 Neural Style Transfer
Recent image generation and style transfer algorithms (Liu et al., 2021; Heitz et al., 2021; Jing et al., 2020; Yoo et al., 2019; Risser et al., 2017; Gatys et al., 2017; Chen et al., 2021a; Wang et al., 2020) use deep learning, such as CNNs, to measure the texture similarity between two images in terms of the distance between their neural activations. Specifically, feature maps of a pre-trained classifier such as VGG-19 (Simonyan & Zisserman, 2014) are extracted and low-order moments of their distribution (Kessy et al., 2018; Huang & Belongie, 2017; Sheng et al., 2018) are used as texture descriptors. By matching such feature statistics, CNN-based techniques have shown promising results transferring texture between arbitrary images, significantly improving over classic texture synthesis approaches (Portilla & Simoncelli, 2000; Zhu et al., 2000; Heeger & Bergen, 1995).
A large body of work in CNN-based style transfer focuses on artistic applications such as reproducing an artwork style and imposing it over a scene of interest. These methods adopt either (i) an iterative optimization approach (Heitz et al., 2021; Risser et al., 2017; Gatys et al., 2016; 2017; Li et al., 2017a), where an initial guess is gradually transformed to depict a style of interest or (ii) an encoding-decoding strategy (Liu et al., 2021; Jing et al., 2020; Li et al., 2017b; Chen et al., 2021a; Wang et al., 2020), where one or more CNN-based image generators are trained to impose a target texture over the scene of interest in a single forward pass. While selecting one of these implies a trade-off between synthesis quality and computational cost, in most cases the generated stylization shows an unnatural appearance, i.e., it does not capture most of the texture details of a style image.
Alternative techniques show impressive style generalization with much simpler transformations. In their seminal work, Dumoulin et al. (2017) propose a style transfer technique that extracts a low-dimensional representation from an arbitrary style image. This representation is then used to predict scaling and shifting values that are used to normalize the feature maps of a content image. The authors empirically show that, by learning such a contracted representation of style images, their method can represent styles from a wide range of unseen domains, leading to a faithful stylization.
Other works have applied style transfer in context of supervised learning, to produce additional labelled data (Zheng et al., 2019), remove model biases (Geirhos et al., 2018) and improve robustness against adversarial attacks (Hong et al., 2021). To our knowledge, this is the first work using neural style transfer to learn more performant self-supervised representations, without labelled data.
3 Preliminaries
3.1 Self-Supervised Learning
Our proposed work deals with data augmentation in SSL. More precisely, we propose a novel preprocessing operation to enhance representation learning. Before introducing our work, we review the fundamentals of SSL by revisiting a well established technique: SimCLR.
A Simple Framework for Contrastive Learning. SimCLR learns compressed representations by maximizing the agreement between differently augmented views of the same data example in a latent space, while minimizing the similarity to different examples. Potential augmentations include random cropping, flipping, color jitter, blurring, and solarization. By maximizing the similarity of augmented training samples, the network learns to create robust representations that are invariant to simple distortions that could occur in the real world, and which should not affect the semantic content of an image. By simultaneously contrasting against other examples, the representations learn to extract information that is unique to each example. The combination of the two leads to the extraction of useful semantic information from the data.
Given a batch of $N$ input images $\{{\bm{i}}_{k}\}_{k=1}^{N}$, $2N$ augmented samples are generated by applying distinct transformations to each image. These transformations correspond to the same data augmentation pipeline. Let $R$ correspond to all possible augmentations. Then, positive pairs correspond to augmented views of the same input sample, and negative pairs correspond to views coming from different input images. Based on this, the $2N$ augmented samples $\{\tilde{{\bm{i}}}_{l}\}_{l=1}^{2N}$ can be organized so that indices $l=2k-1$ and $l=2k$ correspond to views of the same input sample
$\tilde{{\bm{i}}}_{2k-1}=\ r({\bm{i}}_{k}),\quad\tilde{{\bm{i}}}_{2k}=\ \hat{r}({\bm{i}}_{k}),\quad\hat{r},r\sim R$ (1)
Once augmented views are obtained, a representation is computed using an encoder (typically a CNN-based model). The representations are then fed to a projection head which further compresses the representation into a lower-dimensional manifold where different views of the same image are close together and those from different images are far apart. Let $h$ and $g$ be the encoder (e.g. a ResNet-50 backbone) and projection head (e.g. a multilayer perceptron), respectively. Then, embeddings are obtained for each augmented sample
${\bm{z}}_{l}=g\circ h(\tilde{{\bm{i}}}_{l})$ (2)
SimCLR relies on the normalized temperature-scaled cross entropy loss (NT-Xent) to learn how to distinguish between positive pairs of augmented samples. First, the cosine similarity distance between every pair of embeddings is computed
$s_{m,n}=$ $\ \frac{\langle{\bm{z}}_{m},{\bm{z}}_{n}\rangle}{\|{\bm{z}}_{m}\|\|{\bm{z}}_{n}\|}$ (3)
The final step is to use a contrastive loss to train the model. The contrastive loss compares the embeddings of positives and encourages them to be as similar as possible in the latent space. Since the loss is normalized, it naturally forces the representations of views from two different images (negatives) to be far away from each other.
$\mathcal{L}=$ $\ \frac{1}{2N}\sum_{k=1}^{N}\big{[}l(2k-1,2k)+l(2k,2k-1)\big{]}$ (4)
$l(m,n)=$ $\ -\log\bigg{(}\frac{\exp(s_{m,n}/\tau)}{\sum_{l=1}^{2N}\mathbbm{1}_{m\neq n}\exp(s_{m,l}/\tau)}\bigg{)}$ (5)
where $\tau\in\mathbb{R}_{++}$ is the temperature factor and $\mathbbm{1}$ the indicator function. While SimCLR is a simple framework, it pushed the state-of-the-art significantly on a wide range of downstream tasks including image classification, object detection, and semantic segmentation. Follow up works to SimCLR such as Grill et al. (2020); Chen & He (2021); Caron et al. (2020; 2021); Assran et al. (2022); Zbontar et al. (2021); Chen et al. (2020b; 2021b); Bardes et al. (2021) have proposed modifications to this setup which attempt to further improve the downstream task performance.
3.2 Neural Style Transfer
Neural Style Transfer algorithms aim to preserve the content of an image while imposing the style of another. These approaches assume that the statistics of shallower layers of a trained CNN encode style, while deeper layers encode content. Thus, early Style transfer algorithms worked by passing the pair of content and style images to a trained image encoder and optimizing the content image to produce activations which had similar statistics to the style image at shallower layers while keeping the deeper layer activations unchanged (Simonyan & Zisserman, 2014).
Once semantic and texture features are extracted, a stylized image with the semantic attributes of the content image and the texture properties of the style image is generated. The two main approaches are image-optimization based and model-based methods. In what follows, we introduce a model-based Style Transfer technique adopted by our proposed method due on its generalization and efficiency properties. Jing et al. (2019) provide an in-depth study of image stylization techniques.
Fast Style Transfer. Dumoulin et al. (2017) proposed an arbitrary style transfer method with remarkable generalization properties. Their algorithm, Fast Style Transfer, has been shown to accurately represent unseen artistic styles by training a model to predict first and second moments of latent image representations at multiple scales. Such moments are used as arguments of a special form of instance normalization, denominated conditional instance normalization (CIN), to impose style over arbitrary input images. Below we formally introduce the Fast Style Transfer algorithm.
Given a content image ${\bm{i}}_{c}\in\mathbb{R}^{C\times H_{c}\times W_{c}}$ and a style image ${\bm{i}}_{s}\in\mathbb{R}^{C\times H_{s}\times W_{s}}$, the produced stylized image ${\bm{i}}_{\text{cs}}$ corresponds to
${\bm{i}}_{\text{cs}}=$ $\ \mathcal{T}({\bm{i}}_{c},{\bm{z}}_{s})\in\mathbb{R}^{C\times H_{c}\times W_{c}}$ (6)
where $\mathcal{T}$ is a stylization network and ${\bm{z}}_{s}$ is a feature embedding extracted from the style image using a pre-trained feature extractor denoted as $\mathcal{F}$, e.g., InceptionV3 (Szegedy et al., 2016).
${\bm{z}}_{s}=$ $\ \mathcal{F}({\bm{i}}_{s})\in\mathbb{R}^{D}$ (7)
Here, we assume ${\bm{z}}_{s}$ to be a contracted embedding of the style image ($D\ll CH_{s}W_{s}$). The stylization network $\mathcal{T}$ is comprised by $L$ blocks $\{t_{l}\}_{l=1}^{L}$. These extract high-level features from the content image, align them to the style embedding ${\bm{z}}_{s}$ and map the resulting features to the pixel domain.
The style of ${\bm{i}}_{s}$ encapsulated in ${\bm{z}}_{s}$ is transferred to the content image using CIN. This is applied to a particular set of layers to impose the target texture and color scheme by aligning feature maps at different scales. We define the set of layers where CIN is applied as $\mathcal{A}$. Conditional instance normalization (CIN) consists of an extended form of instance normalization where the target mean and standard deviation are extracted from an arbitrary style representation ${\bm{z}}$. Given an input ${\bm{i}}\in\mathbb{R}^{C\times H\times W}$ and a style representation ${\bm{z}}\in\mathbb{R}^{D}$, CIN is defined as
$\hat{{\bm{i}}}=$ $\ \text{CIN}({\bm{i}},{\bm{z}})$ (8)
$\hat{{\bm{i}}}^{(k)}=$ $\ \gamma^{(k)}({\bm{z}})\bigg{(}\frac{{\bm{i}}^{(k)}-\mathbb{E}[{\bm{i}}^{(k)}]}{\sigma({\bm{i}}^{(k)})}\bigg{)}+\lambda^{(k)}({\bm{z}}),\ k\in\{1,\dots,C\}$ (9)
where ${\bm{i}}^{(k)}$ corresponds to the $k$-th input channel, and the sample mean and standard deviation are computed along its spatial support. Here, $\gamma^{(k)},\lambda^{(k)}:\mathbb{R}^{D}\mapsto\mathbb{R}$ are trainable functions that predict scaling and offset values from the latent representation ${\bm{z}}$ for the $k$-th input channel. Following this, the layers in the stylization network are characterized by
$\bm{\phi}_{c,l}=$$\ \begin{cases}\text{CIN}_{l}\big{(}t_{l}(\bm{\phi}_{c,l-1}),{\bm{z}}_{s}\big{)},&l\in\mathcal{A}\\ t_{l}(\bm{\phi}_{c,l-1}),&l\notin\mathcal{A}\end{cases}$ (10)
where the network input corresponds to $\bm{\phi}_{c,0}={\bm{i}}_{c}$. We use the subscript $l$ in the CIN operation to indicate that each layer has its own $\gamma$ and $\lambda$ functions to normalize feature maps independently.
Our method uses the Fast Style Transfer algorithm. Given its generalization properties and low-dimensional style representations, it is a good match for our framework, where style representations from multiple domains must be extracted, manipulated and transferred in an efficient manner.
4 Proposed Method: SASSL
Recent work in self-supervised learning has shown the importance of diverse data augmentation to achieve state-of-the-art performance across downstream tasks. However, few approaches focus on transformations that preserve the semantic information of pre-training data, though this is essential for strong self-supervised representations. To tackle this limitation, we propose SASSL, a novel data-augmentation strategy based on neural style transfer that decouples semantic from appearance attributes, enforcing transformations that operate strictly on the input color and texture.
Style Transfer as data preprocessing. We incorporate style transfer to the default preprocessing pipeline of SSL methods. We note that our method is not specific to a particular SSL method, and can be readily applied with different methods. Figure 1 describes an example of our SASSL data augmentation pipeline, where Style Transfer is applied after random cropping. A raw input image ${\bm{i}}_{0}$ is randomly cropped, producing a view that is taken as the content image ${\bm{i}}_{c}$. Given an arbitrary style image ${\bm{i}}_{s}$ (we discuss the choice of ${\bm{i}}_{s}$ below), the style transfer block generates a stylized image ${\bm{i}}_{\text{cs}}$ by imposing the texture attributes of ${\bm{i}}_{s}$ over ${\bm{i}}_{c}$. Finally, the stylized image ${\bm{i}}_{\text{cs}}$ is passed to the remaining data augmentation blocks to produce an augmented sample ${\bm{i}}_{\text{aug}}$.
As discussed in recent work on self-supervised augmentation (Han et al., 2022; Chen et al., 2020a), adding a strong transformation to a self-supervised method tends to degrade performance. For this reason, it is crucial to control the amount of stylization imposed in the augmentation stage. We do so by introducing three hyperparameters: a stylization probability $p\in[0,1]$, which dictates whether an image is stylized or not, a blending factor $\alpha\in[0,1]$ to combine content ${\bm{z}}_{c}$ and style ${\bm{z}}_{s}$ representations, and an interpolation factor $\beta\in[0,1]$ to combine content ${\bm{i}}_{c}$ and stylized ${\bm{i}}_{\text{cs}}$ images.
Given style representations extracted from content and style reference images ${\bm{z}}_{c}=\mathcal{F}({\bm{i}}_{c})$ and ${\bm{z}}_{s}=\mathcal{F}({\bm{i}}_{s})$, respectively, we obtain an intermediate stylized image $\hat{{\bm{i}}}_{\text{cs}}$ by applying a convex combination on them based on the blending factor $\alpha$.
$\hat{{\bm{i}}}_{\text{cs}}=$ $\ \mathcal{T}({\bm{i}}_{c},\hat{{\bm{z}}})$ (11)
$\hat{{\bm{z}}}=$ $\ (1-\alpha){\bm{z}}_{c}+\alpha{\bm{z}}_{s}$ (12)
Then, the final stylization output is obtained via a convex combination between the intermediate stylized image $\hat{{\bm{i}}}_{cs}$ and the content image ${\bm{i}}_{c}$ based on the interpolation factor $\beta$.
${\bm{i}}_{\text{cs}}=$ $\ (1-\beta){\bm{i}}_{c}+\beta\hat{{\bm{i}}}_{\text{\text{cs}}}$ (13)
Algorithm 1 describes our proposed style transfer data augmentation block. Figure 3 illustrates the effect of the feature blending and image interpolation operations, showcasing their importance to control the stylization effect without degrading the semantic attributes.
Input: ${\bm{i}}_{c},{\bm{i}}_{s},\mathcal{F},\mathcal{T},\alpha_{\text{min}},\alpha_{\text{max}},\beta_{\text{min}},\beta_{\text{max}}$
Output: ${\bm{i}}_{\text{cs}}$
${\bm{z}}_{c}\leftarrow\mathcal{F}({\bm{i}}_{c})$ ;
  # Style representation of content image
${\bm{z}}_{s}\leftarrow\mathcal{F}({\bm{i}}_{s})$ ;
  # Style representation of style image
$\alpha\sim\mathcal{U}(\alpha_{\text{min}},\alpha_{\text{max}})$;
  # Blending factor
$\hat{{\bm{z}}}\leftarrow(1-\alpha){\bm{z}}_{c}+\alpha{\bm{z}}_{s}$;
  # Feature blending
$\hat{{\bm{i}}}_{\text{cs}}\leftarrow\ \mathcal{T}({\bm{i}}_{c},\hat{{\bm{z}}})$ $\beta\sim\mathcal{U}(\beta_{\text{min}},\beta_{\text{max}})$;
  # Interpolation factor
${\bm{i}}_{\text{cs}}\leftarrow(1-\beta){\bm{i}}_{c}+\beta\hat{{\bm{i}}}_{\text{cs}}$;
  # Stylized image
Algorithm 1 Style transfer augmentation block
SASSL operates over minibatches, allowing efficient data preprocessing. Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ and ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{s}\times W_{s}}$ be content and style minibatches, respectively, comprised by $B$ images ${\bm{I}}_{c}^{(b)}$ and ${\bm{I}}_{s}^{(b)},b\in\{1,\dots,B\}$. Then, the stylized minibatch ${\bm{I}}_{\text{cs}}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying style transfer between a sample from the content batch and a sample from the style batch, given an arbitrary selection criterion. In what follows, we propose two alternatives for selecting style images to balance between augmentation diversity and efficiency.
Diversifying style references. In contrast to traditional data augmentation, style transfer can leverage a second dataset to extract style references. This opens the possibility of selecting style images from different domains, diversifying the transformations applied to the pre-training dataset. SASSL relies on two approaches for sampling style references: external and in-batch stylization.
External stylization consists on pre-computing representations of an arbitrary style dataset and sampling from them during pre-training. This allows controlling the styles to impose on the augmented views while reducing the computational overhead of Style Transfer. Under this configuration, the Style Transfer block receives a content minibatch along with a minibatch of pre-computed style representations extracted from an arbitrary style dataset, and generates a minibatch of stylized images.
On the other hand, in-batch stylization uses the styles depicted in the content dataset itself by using other images of the content minibatch as style references. This is of particular interest for large-scale pre-training datasets covering multiple image categories and thus textures (e.g. ImageNet). So, enabling the use of a single dataset for both pre-training and stylization is a valid alternative.
Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ be a pre-training minibatch comprised by $B$ images and ${\bm{I}}_{c}^{(b)}\in\mathbb{R}^{C\times H_{c}\times W_{c}},b\in\{1,\dots,B\}$ the image in the $b$-th index. Then, samples from the same minibatch can be used as style references by associating pairs of images in a circular fashion. More precisely, a style minibatch ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying a circular shift on the content minibatch indices
${\bm{I}}_{s}^{(b)}=$ $\ {\bm{I}}_{c}^{((b-b_{0})\ \text{mod}\ B)}$ (14)
where mod denotes the modulo operation and $b_{0}$ is an arbitrary offset. Figure 3 shows ImageNet samples stylized using pre-computed features from Painter by Numbers via external stylization, as well as using other ImageNet samples taken from the content minibatch via in-batch stylization.
Figure 2: Feature blending and image interpolation. A fine-grained control over the final stylization is obtained by introducing interpolation factors $\alpha$ and $\beta$ to operate in the feature space and pixel domain, respectively.
Content Image Style (In-batch) Style (External) Stylized (In-batch) Stylized (External) Refer to caption
Figure 2: Feature blending and image interpolation. A fine-grained control over the final stylization is obtained by introducing interpolation factors $\alpha$ and $\beta$ to operate in the feature space and pixel domain, respectively.
Figure 3: Style transfer examples. Stylized images generated using style references from the same domain (in-batch) as well as from other domain (external). Stylization obtained using a blending factor $\alpha=0.5$.
5 Experiments
5.1 Downstream Task Performance
We compare the downstream ImageNet classification accuracy of SSL models learned using SASSL against a MoCo v2 (Chen et al., 2020b) baseline learned using the default data augmentation. Note that MoCo v2 and SimCLR are extremely similar (same loss, backbone architecture, and default augmentations), and differ mainly by MoCo’s use of a momentum encoder.
Pre-training settings. Our pre-training setup is similar to the canonical SSL setup used to pre-train SimCLR and BYOL. We use the same loss, architecture, optimizer, and learning rate schedule as MoCo v2 for fair comparison. We pre-train a ResNet-50 encoder on ImageNet for $1,000$ epochs using our proposed method. Once the backbone is pre-trained, to measure downstream accuracy, we add a linear classification head on top of the backbone and train in a supervised fashion on ImageNet.
SASSL applies Style Transfer only to the left view during pre-training, i.e., no changes in augmentation are applied to the right view. Style transfer is applied with a probability $p=0.8$ using blending and interpolation factors drawn from a uniform distribution $\alpha,\beta\sim\mathcal{U}(0.1,0.3)$. We found that this modest application of style transfer best complimented the existing augmentations, avoiding overly-strong augmentations that can reduce performance (Han et al., 2022; Chen et al., 2020a).
Results. Table 1 compares the downstream classification accuracy obtained by our SASSL augmentation approach on MoCo v2 using external stylization from the Painter by Numbers dataset (Kan, 2016). Results indicate our proposed augmentation improves downstream task performance by $2.09\%$ top-1 accuracy. This highlights the value of Style Transfer augmentation in self-supervised training, where downstream task performance significantly boosts by incorporating transformations that decouple content and style. We also report results with in-batch stylization in Section 5.3.
Please refer to Section A.10 for additional ImageNet results showing SASSL’s performance on larger models (ResNet-50 x4 pre-trained via MoCo v2) and different architectures (ViT-B$/16$ pre-trained via MoCo v3). We experimentally show that SASSL also improves downstream task performance under these scenarios, highlighting its potential across different model complexities and architectures.
Table 1: Downstream classification performance on ImageNet. Classification accuracy of MoCo v2 pre-trained using our proposed style transfer augmentation approach (SASSL). Mean and standard deviation reported over five randomly initialized models.
[…]
MoCo v2 (default) $72.55\pm 0.67$ $91.19\pm 0.34$
SASSL $+$ MoCo v2 (Ours) $\bf 74.64\pm 0.43$ $\bf 91.68\pm 0.36$
Table 2: Transfer learning performance. Downstream top-1 classification accuracy of MoCo v2 pre-trained on ImageNet using SASSL. SASSL generates specialized representations that improve transfer learning performance, both in linear probing and fine-tuning. Mean and standard deviation reported over five randomly initialized models.
Target Dataset
ImageNet ImageNet (1%) iNat21 Retinopathy DTD
Style Dataset Linear Probing
None (Baseline) $72.55\pm 0.67$ $53.23\pm 0.45$ $41.33\pm 0.2$ $\bf 75.88\pm 0.12$ $72.68\pm 0.7$
ImageNet (Ours) $74.07\pm 0.46$ $56.87\pm 0.43$ $45.01\pm 0.04$ $75.75\pm 0.1$ $73.69\pm 1.22$
iNat21 (Ours) $74.28\pm 0.38$ $56.76\pm 0.23$ $44.70\pm 0.37$ $75.75\pm 0.17$ $72.75\pm 1.01$
Retinopathy (Ours) $74.02\pm 0.61$ $\bf 56.99\pm 0.26$ $44.9\pm 0.16$ $75.78\pm 0.08$ $73.73\pm 0.57$
DTD (Ours) $74.32\pm 0.37$ $56.77\pm 0.36$ $\bf 45.08\pm 0.31$ $75.76\pm 0.11$ $\bf 74.41\pm 1.39$
PBN (Ours) $\bf 74.64\pm 0.43$ $56.9\pm 0.18$ $45.02\pm 0.14$ $75.79\pm 0.07$ $72.77\pm 0.77$
Fine-tuning
None (Baseline) $74.89\pm 0.67$ $51.61\pm 0.13$ $77.92\pm 0.14$ $78.89\pm 0.2$ $71.54\pm 0.43$
ImageNet (Ours) $75.52\pm 0.23$ $51.74\pm 0.14$ $79.21\pm 0.07$ $79.64\pm 0.16$ $\bf 72.31\pm 1.85$
iNat21 (Ours) $\bf 75.58\pm 0.47$ $\bf 51.86\pm 0.3$ $79.19\pm 0.12$ $79.6\pm 0.23$ $71.35\pm 1.58$
Retinopathy (Ours) $75.52\pm 0.64$ $51.76\pm 0.26$ $79.23\pm 0.05$ $79.63\pm 0.13$ $72.07\pm 1.61$
DTD (Ours) $75.24\pm 0.65$ $51.73\pm 0.19$ $\bf 79.24\pm 0.08$ $\bf 79.7\pm 0.15$ $70.59\pm 1.42$
PBN (Ours) $75.05\pm 0.69$ $51.85\pm 0.16$ $79.2\pm 0.1$ $79.63\pm 0.13$ $71.35\pm 0.96$
5.2 Transfer Learning Performance
To understand robustness and generalization of representations learned using our approach, we study transfer learning performance across five tasks. By incorporating Style Transfer, we hypothesize that learned representations become invariant to changes in appearance (e.g. color and texture). This forces the feature extraction to rely exclusively on semantic attributes. As a result, the learned representations may become more robust to domain shifts, improving downstream task performance across datasets. We empirically show this by evaluating the transfer learning performance of representations trained using SASSL.
Downstream training settings. We compare the transfer learning performance of ResNet-50 pre-trained via MoCo v2 using SASSL against a vanilla MoCo v2 approach using default augmentations. Models are pre-trained on ImageNet and transferred to four target datasets: ImageNet 1% subset (Chen et al., 2020a), iNaturalist ‘21 (iNat21) (iNaturalist, 2021), Diabetic Retinopathy Detection (Retinopathy) (Kaggle & EyePacs, 2015), and Describable Textures Dataset (DTD) (Cimpoi et al., 2014). We also evaluate downstream performance on ImageNet, giving a total of five target datasets.
To have a clear idea of the effect of the style dataset in the SASSL pipeline, we pre-train five ResNet-50 backbones, each using a different style source. We cover the following style datasets: ImageNet, iNat21, Retinopathy, DTD, and Painter by Numbers (PBN). More precisely, we transfer five models, each pre-trained on a different style, to each of five target datasets. We also transfer a model pre-trained using the default augmentation pipeline. This leads to $30$ different transfer learning scenarios used to better understand the effect of various styles on different image domains.
Transfer learning performance is evaluated in terms of top-1 classification accuracy on linear probing and fine-tuning. All models were pre-trained as described in Section 5.1. We report mean and standard deviation across five model initializations. Please refer to Section A.4 for linear probing and fine-tuning training and testing settings.
Results. Table 2 shows the top-1 classification accuracy obtained via transfer learning. For linear probing, SASSL significantly improves the average performance on four out of five target datasets by up to $3.75\%$ top-1 classification accuracy. For Retinopathy, our augmentation method obtains on-par linear probing performance with respect to the default MoCo v2 model.
Similarly, for fine-tuning, all models trained via SASSL outperform the baseline. Results show the average top-1 classification accuracy improves by up to $1.32\%$. This suggests SASSL generalizes across datasets, spanning from texture (DTD) to the medical images (Retinopathy). Note that, for a fair comparison, we do not perform hyperparameter tuning on the default augmentations.
Table 3: Ablation Study. Linear probing accuracy for representations trained under different configurations. Mean and standard deviation reported over five randomly initialized models.
[…]
MoCo v2 (default) -- -- 72.55±.67plus-or-minus72.55.6772.$5$67 $91.19\pm 0.34$
SASSL $+$ MoCo v2 (Ours) Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation $\beta=1$ ImageNet (in-batch) $70.87\pm 0.72$ $89.33\pm 0.34$
Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation: $\beta\in[0.1,0.3]$ ImageNet (in-batch) $74.07\pm 0.62$ $91.58\pm 0.18$
Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation: $\beta\in[0.1,0.3]$ PBN (external) $\bf 74.64\pm 0.43$ $\bf 91.68\pm 0.36$
5.3 Ablation Study
To better understand the impact of each proposed component, we conduct an ablation study over SASSL’s hyperparameters. Similar to previous experiments, we quantify our uncertainty by reporting the mean and variance obtained by five independent trials.
Downstream training settings. For the ablation study, we cover four scenarios: (i) default MoCo v2, (ii) SASSL $+$ MoCo v2 using only in-batch representation blending, i.e., without interpolation ($\beta=1$), (iii) SASSL $+$ MoCo v2 using in-batch representation blending and pixel interpolation, and (iv) SASSL $+$ MoCo v2 using all its attributes (representation blending, pixel interpolation and an external style dataset).
Results. Table 3 shows our ablation study on MoCo v2. It highlights the importance of controlling the amount of stylization using both representation blending and image interpolation. Without image interpolation, using style transfer as augmentation degrades the downstream classification performance by more than $1.5\%$ top-1 accuracy.
On the other hand, by balancing the amount of stylization via blending and interpolation, SASSL boosts performance by more than $1.5\%$. This is a significant improvement for the challenging ImageNet scenario. Finally, by incorporating an external dataset such as Painters by Numbers, we further improve downstream task performance by almost $2.1\%$ top-1 accuracy. This shows the importance of diverse style references and their effect on downstream tasks.
Please refer to Section A.7 for additional ablation studies on the effect of the style latent representation ${\bm{z}}_{s}$ and the number of stylized layers on the downstream accuracy.
6 Conclusion
We propose SASSL, a novel data augmentation approach based on Neural Style Transfer that exclusively transforms the style of training samples, diversifying data augmentation during pre-training while preserving semantic attributes. We empirically show our approach outperforms well-established methods such as MoCo v2 by more than $2\%$ top-1 classification accuracy on the challenging ImageNet dataset. Importantly, SASSL also improves the transfer capabilities of learned representations, enhancing linear probing and fine-tuning classification performance across image domains by more than $3\%$ and $1\%$ top-1 accuracy, respectively. Future work includes extending our technique to different SSL methods and encoder architectures (e.g. ViTs), which can readily be combined with our method.
7 Acknowledgements
The authors would like to thank Arash Afkanpour and Luyang Liu for their insightful comments and feedback.
Appendix A Appendix
The appendix is organized as follows
  • Sec. A.1 shows the details of the target and style datasets used in our downstream evaluations.
  • Sec. A.2 provides insight on how to select an external style dataset by analyzing their style representations.
  • Sec. A.3 includes additional information on the pre-training settings used in our experiments.
  • Sec. A.4 includes detailed information on the downstream task settings used in our experiments.
  • Sec. A.5 shows additional experiments on downstream performance using SimCLR and BYOL.
  • Sec. A.6 reports few-shot classification performance of MoCo v2 equipped with SASSL.
  • Sec. A.7 includes ablation studies to better understand the effect of each component of Neural Style Transfer (NST) in the downstream performance.
  • Sec. A.8 covers the computational requirements of our proposed method.
  • Sec. A.9 covers recent work on semantic-aware data augmentation and NST, and discusses their differences with our proposed method.
  • Sec. A.10 shows additional experiments on downstream performance using a more complex representation model (ResNet-50 x4) and an alternative representation architecture (ViT-B$/16$).
A.1 Target and Style Datasets
We provide the details of the image datasets used in our experiments. Table 4 covers both target and style datasets, including their size and splits.
Table 4: Target and Style Datasets. Additional details on data split and number of samples of the image datasets used in our experiments.
Dataset Task Train Split Val. Split Test Split
ImageNet Pre-training, Target, Style $1,281,167$- $50,000$
ImageNet $1\%$ Target $12,811$- $50,000$
i_Naturalist 2021 Target, Style $2,686,843$- $500,000$
Diabetic Retinopathy Detection Target, Style $35,126$ $10,906$ $42,670$
Describable Textures Dataset Target, Style $1,880$ $1,880$ $1,880$
Painter by Numbers Style $79,433$- $23,817$
A.2 Style Dataset Selection
Our transfer learning results in Section 5 demonstrate that SASSL achieves improved or comparable downstream performance across multiple datasets by incorporating NST as data augmentation. This raises an important question: how can we select an external style dataset to ensure downstream accuracy improvement? Here, we delve into the similarity between datasets in terms of their styles and establish its connection to the performance improvement gained by using them as external style references.
We focus on the linear-probing scenario as it freezes the representation model, forcing the classification head to rely on the representations learned during pre-training (rather than updating them as is the case with fine-tuning) to distinguish between the target categories.
As an example, in our linear probing experiments presented in Table 2, when Diabetic Retinopathy is used as the target dataset, the downstream accuracy achieved via SASSL + MoCo v2 is comparable to that of the default MoCo v2 algorithm, meaning there is no improvement in performance. We hypothesize that this is because the style representations of Diabetic Retinopathy are significantly different from those of the pre-training (ImageNet) and style datasets. Therefore, learning representations that are invariant to such a distinct set of styles does not contribute to distinguishing between target classes.
To support our hypothesis, we visualize the relationship between style representations using low-dimensional embeddings generated via t-SNE (Van der Maaten & Hinton, 2008) to capture the similarity between styles of different datasets. Style representations of each dataset, corresponding to vectors of length $100$, are obtained using an InceptionV3 feature extractor, as done by the FastSt algorithm. Next, we randomly select 1,800 style representations from each class and compute their two-dimensional embeddings using a perplexity of $30$, early exaggeration of $12$, and initializing the dimensionality reduction using PCA. We compute embeddings using 2,048 iterations.
Figure 4 depicts the low-dimensional embeddings obtained via t-SNE from all datasets used in our transfer learning experiments. The low-dimensional representation shows that the style representations from Diabetic Retinopathy are significantly distinct from those of the rest of datasets, including ImageNet. This aligns with our hypothesis, suggesting that SASSL improves transfer learning when the pre-training and style references are similar to those of the target dataset. From the perspective of t-SNE embeddings, this implies that pre-training and style datasets must have a good overlap with the target dataset for SASSL to improve downstream performance.
Figure 4: t-SNE visualization of style representations. Two-dimensional embeddings of the style representations of different datasets, extracted by the Fast Style Transfer method. Style embeddings of the Diabetic Retinopathy dataset (marked in yellow) form clusters that do not overlap with the rest of datasets, while embeddings from the remaining datasets are close to each other.
A.3 Additional Experimental Details
All of our experiments were implemented in Tensorflow (Abadi et al., 2016). All our models were pre-trained on $64$ TPUs using a batch size of $4,096$. Both the MoCo v2 models pre-trained using the default augmentation and our proposed SASSL approach do not use a dictionary queue.
We used a ResNet-50 (He et al., 2016) as our representation backbone. For our projection head, we used a Multilayer perceptron with $4,096$ hidden features, and an output dimensionality of $256$. Our left tower used a prediction network with the same architecture as the projector, similar to the setup used in BYOL (Grill et al., 2020). Our right tower was a momentum encoder, having the same encoder and projector as the left tower, but whose parameters were an exponential moving average of the corresponding parameters in the left tower and were not trained via gradient descent. Similar to previous works, we used a momentum which started at $0.996$, and which followed a cosine decay schedule ending at 1.0. For the pretraining loss, we used the InfoNCE loss with a temperature of 0.1, similar to what was done in both MoCo v2 and SimCLR.
For our pretraining augmentations, we followed the setup used in BYOL (Grill et al., 2020). The operations used and their hyperparameters (in order of application) are as follows:
  1. 1.
    Random cropping and rescaling to $224\times 224$ with the area chosen randomly between $0.08$ and $1.0$ of the original image and with a logarithmically distributed axis ratio between $3/4$ and $4/3$. This was applied with a probability of 1.0 since it was necessary to get a fixed image shape.
  2. 2.
    Random horizontal flipping with a probability of $0.5$ that it will be applied.
  3. 3.
    Random color jitter. Color jitter consists of $4$ independent transformations, each of which is applied in a random order with randomly chosen values. This transform is described in greater detail in Chen et al. (2020a) and Grill et al. (2020). We used the same configuration as used in those papers.
  4. 4.
    Random grayscaling with a probability of 0.2 that it will be applied.
  5. 5.
    Random blurring with a kernel width distributed randomly between 0.1 and 2.0 pixels. Similar to Grill et al. (2020), we used a probability of $1.0$ for the left view, and a probability of $0.1$ for the right view.
  6. 6.
    Random solarization, which was only applied to the right view with a threshold of $0.5$ and a probability of $0.2$ that it will be applied.
When using SASSL, we applied style transfer after random cropping and before random horizontal flipping. Our default hyperparameters for SASSL were blending and interpolation factors drawn randomly from a uniform distribution between $0.1$ and $0.3$, and a probability of $0.8$ that style transfer will be applied.
For optimization, we used the LARS optimizer (You et al., 2017) with a cosine decayed learning rate warmed up to $4.8$ over the course of the first $10$ epochs. Similar to previous works, we used a trust coefficient of $0.001$, exempted biases and batchnorm parameters from layer adaptation and weight decay, and used a weight decay of $1.5\times 10^{-6}$.
A.4 Downstream Training and Testing Settings
For performance evaluation on downstream tasks, all our models were trained on $64$ TPUs, but using a batch size of $1,024$. In this section, we provide additional details of the downstream training configuration used in our experiments. These cover data augmentation, optimizer and scheduler settings for both linear probing and fine-tuning scenarios.
Linear Probing Settings. We base our linear probing settings on those used by well-established SSL methods (Grill et al., 2020; Chen et al., 2020a; Kornblith et al., 2019) with some changes on the optimizer settings. We also adapt the augmentation pipeline based on the target dataset.
In all our linear probing experiments, the optimization method corresponds to SGD with Nesterov momentum using a momentum parameter of $0.9$. We use an initial learning rate of $0.2$ and no weight decay. We use a cosine scheduler with no warmup epochs and a decay factor of $10^{-6}$. Similarly to previous work, for datasets including a validation split, we trained the linear probe on the training and validation splits together, and evaluated on the testing set.
For small target datasets (ImageNet $1\%$, Retinopathy, and DTD), models were trained for $5,000$ iterations using a batch size of $1,024$, which is consistent with the $20,000$ iterations using a batch size of $256$ reported by previous methods. No data augmentation is applied during training. Instead, during both training and testing, images are resized to $224$ pixels along the shorter dimension followed by a $224\times 224$ center crop and then standardized using the ImageNet statistics.
For iNat21, comprised by $2.6$ million training images, we train the linear probe for $90$ epochs. We empirically found that longer training significantly improved the downstream classification performance both for our proposed SASSL pipeline as well as the default augmentation pipeline.
Similarly, for ImageNet, comprised by $1.2$ million training images, we also train the linear probe for $90$ epochs. Additionally, we included random cropping, horizontal flipping and color augmentations (grayscale, solarization and blurring) during training.
Fine-tuning Settings. Our fine-tuning configuration follows the one used for linear-probing. In all cases, We use SGD with Nesterov momentum using a momentum parameter of $0.9$. Training uses an initial learning rate of $0.2$ and no weight decay. We use a cosine scheduler with no warmup epochs and a decay factor of $10^{-6}$. Similarly to previous work, for datasets including a validation split, we fine-tune the model on the training and validation splits together, and evaluate on the testing set.
The number of training iterations and data augmentation depend on the target dataset, and are identical to those used for linear probing. Note that we do not run a hyperparameter sweep for selecting either the weight decay or initial learning rate, i.e., these remain fixed for all experiments.
A.5 Downstream Task Performance on other SSL Methods
We extend the downstream performance experiments reported in Sec. 5.1 by evaluating SASSL’s impact on two other well-established SSL methods: SimCLR and BYOL. We pre-trained a ResNet-50 backbone using each method, both with and without SASSL, and then evaluated the resulting models on the ImageNet dataset using linear probing. In both cases, we employed the default pre-training and linear probing configurations provided by each method. SASSL was implemented with its recommended hyperparameters (blending and interpolation factor between $0.1$ and $0.3$, probability $0.8$, and Painter by Numbers as external style dataset). The pre-training and fine-tuning settings for SimCLR and BYOL adhered to their default configurations.
Table 5 shows the top-1 and top-5 downstream classification accuracy. Results shows that SASSL consistently improves the downstream classification accuracy of both SimCLR and BYOL, suggesting its effectiveness across different SSL techniques.
Table 5: SimCLR and BYOL Downstream classification performance on ImageNet. Classification accuracy of SimCLR and BYOL equipped with our proposed style transfer augmentation approach (SASSL). Top-1 and top-5 accuracy reported on a single random seed.
[…]
SimCLR (default) $68.62$ $88.7$
SASSL $+$ SimCLR (Ours) $\bf 69.58$ $\bf 89.01$
BYOL (default) $74.09$ $91.83$
SASSL $+$ BYOL (Ours) $\bf 75.13$ $\bf 92.12$
Table 6: Few-shot classification performance. One-shot and ten-shot classification performance of representations extracted via SASSL $+$ MoCo v2 pretrained on ImageNet.
[…]
MoCo v2 (default) $19.56$ $45.05$
SASSL $+$ MoCo v2 (Ours) $\bf 20.55$ $\bf 46.73$
A.6 Few-shot Classification
To further demonstrate the enhanced representation learning capabilities of SASSL, we conducted additional experiments on another downstream task: few-shot classification. We evaluated SASSL $+$ MoCo v2 representations against the default MoCo v2 representations in the context of one-shot and ten-shot settings on the challenging ImageNet dataset. Note that we use the official TensorFlow Dataset’s partitions, namely imagenet2012_fewshot/1shot and imagenet2012_fewshot/10shot.
Table 6 presents the classification accuracy in both scenarios. Results reveal that SASSL boosts few-shot classification top-1 accuracy by over $1\%$ in both cases. This aligns with our classification experiments, suggesting that SASSL promotes more general image representations.
A.7 Additional Ablation Studies
Effect of the Style Representation in Downstream Performance. We conduct additional experiments to better understand the effect of the style representation ${\bm{z}}_{s}$ in the downstream task performance of models pre-trained via SASSL. Specifically, we replace the style latent code, originally taken from a style image, by (i) i.i.d. Gaussian noise ${\bm{z}}_{s}\sim{\mathcal{N}}(\bm{\mu},\bm{\Sigma})$, and (ii) the style representation of the content image ${\bm{z}}_{s}={\bm{z}}_{c}$. Note that latter case is equivalent to using the stylization model ${\mathcal{T}}$ as an autoencoder, since no style is imposed over the feature maps of the content image.
In both cases, we pre-train and linearly probe a ResNet-50 backbone on ImageNet using MoCo v2 equipped with SASSL. All training and downstream task settings follow our default configurations, as covered in the Appendix Section A.3 and A.4. We also use the default blending and interpolation factors $\alpha,\beta\sim{\mathcal{U}}(0.1,0.3)$.
Table 7 shows the linear probing performance obtained by the two scenarios of interest. As reference, we also include the performance of MoCo v2 with the default data augmentation, as well as MoCo v2 via SASSL using an external style dataset (Painter by Numbers). Results show that using noise as style representation boosts top-1 accuracy by $0.24\%$ with respect to the default data augmentation, while using the content as style reference improves performance by $0.8\%$. This implies that using noise as style representation hinders performance with respect to just encoding the input image. On the other hand, using an external style dataset boosts up performance by $2.4\%$, which is a significantly larger improvement over the two scenarios of interest.
Results suggest that the style reference has a strong effect on the downstream performance of the pre-trained models. Either by replacing the latent representation of a style image by noise or removing the style alignment process and keeping the compression induced by the Stylization network ${\mathcal{T}}$ (by forcing ${\bm{z}}_{s}={\bm{z}}_{c}$), the improvement provided by our proposed data augmentation is significantly smaller than that obtained with our full technique using external style images.
Effect of the Number of Stylized Layers in Downstream Accuracy. we conduct an ablation study to investigate the impact of the number of layers stylized during pre-training. We examine three scenarios: (i) stylizing the first two residual blocks (four layers corresponding to Residual blocks 1 and 2), (ii) stylizing the first four residual blocks (eight layers corresponding to Residual blocks 1 to 4), and (iii) stylizing all residual layers (ten layers corresponding to Residual layers 1 to 5).
For each scenario, we pre-trained and linearly probed a ResNet-50 on ImageNet using SASSL $+$ MoCo v2 with its optimal configuration (blending and interpolation between $0.1$ and $0.3$, using Painter by Numbers as the external style dataset). To provide a clear comparison, we compared the classification accuracy of these three models to the model trained without feature alignment (using the stylization network as an autoencoder), and our complete SASSL $+$ MoCo v2 model (stylizing all layers). Note that the latter stylizes all residual and upsampling blocks (a total of thirteen layers).
The results indicate a gradual improvement in accuracy as the number of stylized layers increases. Stylizing only the first four layers yielded no significant improvements, resulting in approximately the same accuracy as the model pre-trained without feature alignment. Stylizing the first eight layers enhanced top-1 classification accuracy by $0.34\%$ over the model with no feature alignment. Finally, stylizing the first ten layers improved accuracy by $0.52\%$. These findings suggest that deeper layers have a more pronounced impact on accuracy improvement. Additionally, the model pre-trained with full stylization achieved a $1.61\%$ improvement over the model with no feature alignment, indicating that the stylization occurring in the upsampling layers plays a dominant role in the gains provided by SASSL.
This study suggests that significant improvements can be achieved by solely stylizing deeper layers, potentially reducing the computational demands of SASSL while still enabling the extraction of robust image representations.
Table 7: Effect of the Style Representation in Downstream Performance. Different style representation settings are analyzed when pretraining a ResNet-50 backbone via MoCo v2. Performance reported on a single random seed.
[…]
Baseline -- -- $72.97$ $90.86$
SASSL (Ours) Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation $\beta\in[0.1,0.3]$ Gaussian Noise ${\bm{z}}_{s}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ $73.21$ $91.17$
${\bm{z}}_{s}={\bm{z}}_{c}$ $73.77$ $91.64$
PBN (external) $\bf 75.38$ $\bf 92.21$
Table 8: Effect of the Number of Stylized layers in Downstream Performance. Downstream classification accuracy of a ResNet-50 backbone pretraining via SASSL $+$ MoCo v2, where NST is applied to a different number of layers. Performance reported on a single random seed.
[…]
Baseline -- -- $72.97$ $90.86$
SASSL (Ours) Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation: $\beta\in[0.1,0.3]$ ${\bm{z}}_{s}={\bm{z}}_{c}$ $73.77$ $91.64$
PBN (external) Stylize first $4$ layers $73.75$ $91.58$
PBN (external) Stylize first $8$ layers $74.09$ $91.76$
PBN (external) Stylize first $10$ layers $74.27$ $91.74$
PBN (external) $\bf 75.38$ $\bf 92.21$
A.8 Computational Requirements
We conducted additional experiments to compare the runtime of our proposed method against the default augmentation pipeline. We measured the throughput (augmented images per second) of SASSL relative to MoCo v2’s data augmentation. The throughput was calculated by averaging 100 independent runs on $128\times 128$-pixel images with a batch size of $2,048$. We also report the relative change, which indicates the percentage decrease in throughput compared to the default data augmentation. All experiments were carried out on a single TPU.
Table 9 summarizes the throughput comparison. SASSL reduces throughput by approximately $20\%$ due to the computational overhead of stylizing large batches, which involves running a forward pass of the NST model. However, empirical evidence shows that SASSL achieves up to a $2\%$ performance improvement. We believe this represents a favorable trade-off between performance and runtime.
Table 9: SASSL runtime. Comparison of the throughput (processed images/second) of SASSL + MoCo v2’s data augmentation pipeline vs. the default MoCo v2’s pipeline.
[…]
MoCo v2 (default) $37.45$-
SASSL $+$ MoCo v2 (Ours) $29.48$ $21.28$
A.9 Comparison to Recent Data Augmentation Work
Neural Style Transfer Augmentation for Self-Supervised Learning. Recent work has explored the use of NST in self-supervised classification models to impose invariances towards improved classifiaction accuracy. Geirhos et al. (2018) studies how CNNs are biased towards texture in the context of supervised learning. Their main hypothesis is that CNN classifiers trained via supervised learning are biased towards texture. This is distinct from our approach, since we focus on preserving semantic information in self-supervised learning by exclusively distorting the style component of pre-training samples. More precisely, we decouple content and style in order to exclusively distort the style to obtain stronger self-supervised image representations. While both works rely on NST, their method uses it to create a stylized dataset (Stylized-ImageNet), whereas our approach is to incorporate style transfer in the augmentation pipeline to create stylized images on-the-fly while controlling their effect via blending and interpolation hyperparameters.
Similarly, Jackson et al. (2019) propose employing NST as a data augmentation technique for supervised learning, where the style reference is drawn from a normal distribution. In contrast, our experiments demonstrate that this approach hinders performance in the context of SSL. SASSL utilizes precomputed style representations from external datasets or samples from the content datasets themselves (in-batch stylization), resulting in improved or on-par downstream performance across style datasets. Furthermore, their approach solely considers the blending of style representations, overlooking the distortions introduced by the stylization network itself. Our work reveals that the compression generated by the stylization network, even without using a style representation, leads to a loss of information about the structure (edges and small image details) of pre-training images, deteriorating downstream performance. This is the rationale behind SASSL’s reliance on both pixel interpolation and feature blending.
Zheng et al. (2019) utilize NST to create additional samples for supervised learning, using only eight style images. It’s important to note that their approach does not incorporate NST as data augmentation; instead, the stylized images are fixed and added to the original training dataset. Additionally, their experiments are limited to a single classification dataset (Caltech). This contrasts sharply with our method, which proposes different stylization strategies (external and in-batch) and ways to control the stylization effect (via ablation and interpolation) to augment data on-the-fly in order to enhance SSL performance. While both methods share the use of NST, their tasks and approaches are distinct.
Semantic-Aware Data Augmentation in Self-Supervised Learning. Purushwalkam & Gupta (2020) suggest an augmentation strategy based on natural (temporal) transformations that occur in videos as an alternative to learning from occluded (cropped) images from object-centric datasets. In contrast, our approach utilizes neural style transfer to decouple images into content and style, allowing us to apply transformations exclusively to the style of pre-training samples while preserving the content information. Their approach relies on datasets generated on video frames to incorporate temporal invariance, while SASSL can be used on standard image datasets, enabling its integration to default augmentation pipelines.
On the other hand, Lee et al. (2021) propose to include an auxiliary loss for SSL methods to capture the difference between augmented views, leading to better downstream performance on datasets where the task relies on information that may have been lost due to invariances introduced by aggressive augmentation. While their method modifies the pre-training loss and requires the use of data augmentation parameters to learn representations, SASSL does not require any additional loss components or auxiliary information from each applied augmentation to function. Our method complements the default data augmentation pipeline with a content-preserving transformation to obtain more general representations.
Finally, Bai et al. (2022) propose an alternative SSL augmentation pipeline to prevent the loss of semantic information in the learned representations by encouraging learning from weakly augmented views and gradually transitioning to strongly augmented views. The method avoids semantic shifts caused by aggressive image transformations by leveraging the tendency of SSL models to memorize clean views during early pre-training stages and by proposing a multi-stage augmentation pipeline to generate weak and strongly augmented views. Differently, our approach does not require additional augmented views or dynamic control over the loss weights during training. SASSL utilizes a content-preserving transformation that only modifies stylistic characteristics (color and texture), leaving the natural image structure and objects unchanged. Our approach incorporates a transformation into the default augmentation pipeline without altering the loss function or instance discrimination strategy.
Table 10: MoCo v2 + SASSL Downstream Performance using a ResNet-50 x4 backbone. Top-1 and Top-5 classification accuracy obtained by pre-training and linear probing on ImageNet. Performance reported on a single random seed.
[…]
MoCo v2 (default) $375$M $77.2$ $93.32$
SASSL $+$ MoCo v2 (Ours) $375$M $\bf 78.21$ $\bf 93.98$
Table 11: MoCo v3 + SASSL Downstream Performance using a ViT-B backbone. Top-1 and Top-5 classification accuracy obtained by pre-training and linear probing on ImageNet. Performance reported on a single random seed.
[…]
MoCo v3 (default) $86$M $75.01$ $92.43$
SASSL $+$ MoCo v3 (Ours) $86$M $\bf 75.51$ $\bf 92.56$
A.10 Downstream Task Performance on Alternative Backbone Architectures
We ran additional experiments to explore the potential of SASSL on models with larger complexity and different architectures. In terms of complexity, we explored the use of ResNet-50 (x4) as representation model, which keeps the same architecture as the ResNet-50 models evaluated in Section 5, but increases the parameters from $23$ million to $375$ million. In terms of architecture, we evaluated the use of ViT-B$/16$, which is comprised by $86$ million parameters.
For comparison purposes, We pre-train and linear probe ResNet-50 (x4) representation models on ImageNet via MoCo v2 with its default augmentation and with SASSL. Both pre-training and downstream settings follow our default configuration, as documented in the Appendix Section A.3 and A.4. Similarly, we pre-train and linear probe ViT-B$/16$ representation models on ImageNet via MoCo v3 with its default augmentation and with SASSL. For the latter case, we updated the blending factor $\alpha$ to be uniformly sampled between $0.1$ and $0.5$ (instead of the default $0.1$ to $0.3$ interval). Note that SASSL does not introduce any additional training parameters.
Tables 10 and 11 report the downstream classification accuracy for ResNet-50 (x4) and ViT-B$/16$, respectively. ResNet-50 (x4) results show SASSL improves top-1 classification accuracy by $1.1\%$. This is consistent with our previous experiments, where ResNet-50 representation models pre-trained with SASSL improved accuracy by $2.1\%$. While still showing a significant improvement, we hypothesize that SASSL’s improvement drop from $2.1\%$ to $1.1\%$ is mainly because the large ResNet-50 (x4)’s model complexity. More precisely, ResNet-50 (x4) is capable of learning better representations than the original ResNet-50, so data augmentation strategies have less effect on its downstream classification accuracy.
On the other hand, ViT-B$/16$ results show SASSL improves downstream top-1 accuracy by $0.5\%$ over the default MoCo v3 model. This suggests that our proposed augmentation is not limited to CNN representation models, but can also be extended to Vision Transformers, although showing a smaller improvement than for CNNs. We hypothetize that this has to do with the way style transfer is applied in our augmentation pipeline. While our stylization model uses CNNs to align content and style features, Vision Transformers extract information via patch-embedding and self-attention layers. So, it is possible that the mismatch between stylization and representation models weakens the stylization effect. An interesting follow-up direction would be to explore data augmentation based on Style Transfer techniques using ViT’s such as the recently proposed by Deng et al. (2022). We speculate that larger improvements may be achieved by matching the stylization and representation models.